19. Investment Criteria. If you insulate your office for S1,000, you will save $100 a year in heating expenses. These avings will last forever a. What is the NPV of the investment when the cost of capital is 8 percent? 10 percent? b. What is the IRR of the investment? What is the payback period on this investment
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B 1 u. Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard AQ190 AO Alignment Number Cells AP AQ AR AS AT AU AV AW AX AY AZ 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 (4 トト MIRR Npu IRR. REPLACEMENTS-HPR GMAM ( EAC MACRS LEASE RATIOCASHBUDGET wacc- BOND EPS EBIT REPLACEMENT STats -NPV BETA LEVERED ☆ perpetual cash flows are given PV OF ALL CASH FLOW = CASHFLOW/REQUIRED RATE NPV AT 8% NPV NPV= NPV = PV OF ALL CASHFLOW - INVESTMENT 100/008-1000 250 NPV AT 10% NPV 100/0.10-1000 NPV = 0 IRR = WHERE NPV = 0 IRR =1096 WE CAN SEE THAT FROM ABOVE ANSWER, WE CAN ALSO CALCULATE AS UNDER NPV = 0 = 100/r-1000 1000 = 100/r r-100/1000-0. 10-10% PV OF ALL CASHFLOW - INVESTMENT erences: E147 福 130% 05:13 09-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard AQ227 AO Font Alignment Number Styles Cells Edting AP AQ AR AS AT AU AV AW AX AY AZ 209 210 211 212 213 214 215 216 217 218 219 220 221 PAYBACK PERIOD AS CASHFLOWS ARE SAME EVERY YEAR, PAYBACK PERIOD INVESTMENT/ANNUAL CFAT PAYBACK PERIOD =1000/100 = 10 YEARS 223 224 225 226 227 /4 MIRR NpV IRR REPLACEMENT S HPR GMAM EAC MACRS LEASE. RATIOCASHBUDGET acc BOND EPS EBIT REPLACEMENT ST BETA LEVERED - erences: E147 福 130% 05:13 09-01-2019